# Federated Sketching LoRA: On-Device Collaborative Fine-Tuning of Large Language Models

### v2

@@ C.2 Further Increasing the Number of Devices
C.3 Varying the Level of Data Heterogeneity
In Table C.3, we investigate the impact of the degree of data heterogeneity on performance. We increase the heterogeneity by decreasing the Dirichlet concentration parameter from $\alpha=1$ to $\alpha=0.1$. The local ranks follow the heavy-tail distribution described in the previous subsection, and all other experimental configurations remain consistent with Appendix C.1. As observed from Table C.3, the performance of all methods degrades as heterogeneity increases. FSLoRA consistently achieves higher accuracy.
Table C.3: Accuracy comparison under different data heterogeneity settings. FSLoRA maintains its advantage over the baselines as the data heterogeneity increases. The number of devices is set to $50$.
@@ C.3 Varying the Level of Data Heterogeneity
Appendix D Experiments on LLaMA-7B
Although our primary focus is on models suitable for on-device deployment, such as RoBERTa and LLaMA-3.2-3B models, we also include experiments on the larger LLaMA-7B model to demonstrate the scalability of FSLoRA in more complex models. Specifically, we fine-tune the LLaMA-7B model on the Wizard, Dolly-15k, and Alpaca datasets and evaluate it on $1444$ MMLU samples (available at: https://github.com/ATP-1010/FederatedLLM). For Wizard and Dolly-15k, we adopt the same heterogeneous data partitioning as [37]. Since the Alpaca dataset lacks a clear task or domain structure, we apply a uniform random partitioning strategy to distribute the data across devices. We tune the q_proj and v_proj modules and set the local LoRA ranks $k_{i}\!=\![64,32,16,16,8,8,4,4,4,4]$ for $10$ devices. The parameter settings are aligned with those in [37].
Table D.1: Performance comparison on LLaMA-7B model.
@@ Appendix D Experiments on LLaMA-7B
As shown in Table D.1, FSLoRA achieves the highest average performance across all three datasets compared to baselines. These results demonstrate FSLoRA’s potential for effective fine-tuning with the large-scale LLaMA-7B model under heterogeneous device settings.
Appendix E Further Experiments
@@
In this section, we provide additional results, including detailed per-task comparisons from the commonsense reasoning benchmark corresponding to Figures LABEL:fig:rank_varying and LABEL:fig:global_rank and the investigation of the integration of communication compression and sketching.
E.1 Further Details on the Ablation Study
@@
Impact of sketching: In Figure 6, we compare the performance of FSLoRA with and without sketching on eight tasks from the commonsense reasoning benchmark using the LLaMA-3.2-3B model. Notably, FSLoRA without sketching is equivalent to the vanilla Federated LoRA. For FSLoRA with sketching, we apply a uniform sketching ratio of $k_{i}/r=0.5$ across all distributed devices. The uploading budget for each device is set to $200$ times the size of the full global LoRA modules at the corresponding rank. It is clear that FSLoRA with sketching consistently outperforms its non-sketched counterpart across these eight tasks, demonstrating the effectiveness of sketching in improving performance.
@@
Figure 6: Comparison of FSLoRA with and without sketching, with an upload budget 200×200$\times$200 × the global LoRA module size at each rank. This is based on the commonsense reasoning benchmark and the LLaMA-3.2-3B model. We observe that the sketching mechanism improves performance across all considered tasks. The average accuracy of the eight tasks is shown in Figure LABEL:fig:rank_varying.
@@
Figure 7: Impact of the rank of global LoRA modules on FSLoRA, given a fixed rank for the updated submatrices at the devices. This is based on the commonsense reasoning benchmark and the LLaMA-3.2-3B model. Overall, FSLoRA demonstrates improved performance as the global rank increases. The average accuracy of the eight tasks is shown in Figure LABEL:fig:global_rank.
@@
Impact of the global rank: In Figure 7, we present the impact of the rank of the global LoRA modules on FSLoRA’s performance across eight tasks from the commonsense reasoning benchmark. We consider four configurations: 1) $r=8,~{}k_{i}/r=1$, 2) $r=16,~{}k_{i}/r=0.5$, 3) $r=32,~{}k_{i}/r=0.25$, and 4) $r=64,~{}k_{i}/r=0.125$. The rank of submatrices updated by the devices at each iteration remains consistent across all configurations (i.e., $k_{i}=8$), ensuring that the communication and computational resources on the client side are kept fixed for all cases. In the ARC-Easy task, performance decreases as the rank increases to $64$, potentially due to overfitting. In general, FSLoRA shows improved performance as the rank increases.
E.2 Integration of Sketching and Top-k Compression
Building on the commonsense reasoning benchmark and the LLaMA-3.2-3B model, we further explore the integration of two orthogonal techniques, sketching and top-k compression. Specifically, we employ top-k compression, a widely used compression approach introduced to LoRA by Kuo et al. [22], in FSLoRA to reduce communication overhead when uploading model updates. In our setup, the compression ratio is fixed at $0.5$ for all methods, while the sketching ratio $k_{i}/r$ varies over $\{0.125,0.25,0.5,1\}$. Notably, FSLoRA with sketching ratio $k_{i}/r=1$ corresponds to the vanilla Federated LoRA (i.e., without sketching). Figure 8 plots testing accuracy versus communication overhead, where the x-axis represents the amount of data uploaded per device (in MB), assuming parameters are stored in float 32 precision. The results clearly show that integrating sketching with top-k compression further improves communication efficiency: methods with lower sketching ratios consistently achieve higher accuracy under the same communication budget, highlighting the potential of FSLoRA for scalable and communication-efficient on-device LLM fine-tuning.
@@ E.2 Integration of Sketching and Top-k Compression
Appendix F Proof of the Theoretical Results
F.1 Preliminaries
Before presenting the proof of the main results, we first introduce some preliminary facts that will be used later.
Lemma F.1.
@@ F.1 Preliminaries
Lemma F.2.
[36, Lemma 2] Suppose a sequence of random matrices $\left\{{\bf P}_{i}\right\}_{i=1}^{N}$ satisfy $\mathbb{E}\left[{\bf P}_{i}\mid{\bf P}_{i-1},\!{\bf P}_{i-2},\ldots,{\bf P}_{1}\right]={\bf 0},\forall i.$ Then,
@@
Lemma F.3.
[20, Lemma 17] For any $a_{0}\geq 0,b\geq 0,c\geq 0,d>0$, there exist a constant $\eta\leq\frac{1}{d}$ such that
@@
Lemma F.4 (Random sketching bounds).
@@
This completes the proof of Lemma F.4. ∎
F.2 Proof of Lemma 3.2
@@
The proof of Proposition F.5 is deferred to Appendix F.5. With this proposition, we are ready to prove Theorem 4.4.
In FSLoRA, the update direction in (4) corresponds to the negative stochastic gradient of $\ell({\bf W}_{0}+{\bf B}{\bf S}{\bf A},\xi)$ with respect to $[{\bf B};{\bf A}]$ for a given sketch ${\bf S}_{i}^{t}$. We have defined ~$\tilde{\ell}({\bf X},\xi;{\bf S})=\ell({\bf W}_{0}+{\bf B}{\bf S}{\bf A},\xi)$. The iterative equation for the proposed FSLoRA algorithm thus can be written as
@@
where ${\bf g}^{t,h}_{i}$ denotes the stochastic gradient $\nabla_{\bf X}\tilde{\ell}({\bf X}^{t,h}_{i},\xi_{i}^{t,h};{\bf S}_{i}^{t})$. Based on the smoothness of $f^{\mathcal{S}}(\bf X)$, i.e., Proposition F.5, we have
@@
where the last inequalities follow Jensen’s inequality and Proposition F.5.
@@
where the inequality follows the fact that $\mathbb{E}[\sum_{h=0}^{H-1}({\bf g}^{t,h}_{i}-\nabla_{\bf X}f_{i}^{\mathcal{S}}({\bf X}_{i}^{t,h}))]=0$ and the independence between devices.
@@
where the equality follows Lemma F.2 and the inequality follows Assumption 4.2. For $\left\|\nabla_{\bf X}f_{i}^{\mathcal{S}}({\bf X}_{i}^{t,h})\right\|^{2}$, utilizing Assumption 4.3 and Proposition F.5, we have
@@
where $T_{3}=\frac{1}{NH}\sum_{i=1}^{N}\frac{r^{2}}{k_{i}^{2}}\sum_{h=0}^{H-1}\mathbb{E}\left\|{\bf X}_{i}^{t,h}-{\bf X}^{t}\right\|^{2}$. Combining (12), (13), and (F.3) yields
@@
Plugging inequality (F.3) into inequality (17) yeilds
@@
Telescoping the above inequality from $t=0$ to $T-1$, we have
@@
where $f^{*}$ denotes the lower bound of $f^{\mathcal{S}}({\bf X})$ and $\widetilde{L}=\kappa L$. This completes the proof of Theorem 4.4.
F.4 Proof of Corollary 4.5
Applying Lemma F.3 to the bound derived in Theorem 4.4 and letting $d=H$, it follows that there exists a learning rate $\gamma\leq\min\{\frac{N}{24\rho(c_{h}+1)H\bar{L}},\frac{1}{8\sqrt{\widetilde{L}L(\rho+1)(c_{h}+1)}H},\frac{1}{H}\}$ such that
@@
This completes the proof of Corollary 4.5.
F.5 Proof of Proposition F.5
i) For illustration, we need to recover $\bf X$ to $[\bf B;{\bf A}]$ in this proof. According to the definition of $\tilde{f}_{i}({\bf X};{\bf S})$ and $f_{i}({\bf B},{\bf A})$, we have
@@
Appendix G Implementation Details
G.1 Details on Hyperparameters
Table G.1: The hyperparameters for RoBERTa & GLUE and LLaMA-3.2-3B & commonsense reasoning benchmarks.
Hyperparameter RoBERTa & GLUE LLaMA-3.2-3B & commonsense reasoning
Dirichlet parameter 0.1 0.1
Batch size 16 16
LoRA dropout rate 0.1 0.1
Learning rate, $\gamma$ 5e-4 3e-4
Communication round, $T$ 200 500
Local iteration number, $H$ 50 20
G.2 Details on Datasets
G.2.1 GLUE Benchmark
[…]
      Input: A single sentence.
[…]
      Input: A single sentence.
    • Output: A label indicating positive or negative sentiment.
[…]
      Input: Two sentences (‘sentence1’ and ‘sentence2’).
    • Output: A label indicating either equivalent or not equivalent.
[…]
      Input: Two questions.
    • Output: A label indicating duplicate or not duplicate.
[…]
    • Output: A label indicating entailment, contradiction, or neutral.
[…]
      Input: A question and a sentence.
  • RTE provides pairs of sentences to see if one implies the other.
[…]
      Input: Two sentences (‘sentence1’ and ‘sentence2’)
G.2.2 Commonsense Reasoning Benchmark
The input template, i.e., prompt format for these datasets is detailed in Table G.2.
Table G.2: The prompt template of the commonsense reasoning datasets [16].
Dataset Input Template

### v3

@@ F.2 Further Increasing the Number of Clients
F.3 Varying the Level of Data Heterogeneity
In Table F.3, we investigate the impact of the degree of data heterogeneity on performance. We increase the heterogeneity by decreasing the Dirichlet concentration parameter from $\alpha=1$ to $\alpha=0.1$. The local ranks follow the heavy-tail distribution described in the previous subsection, and all other experimental configurations remain consistent with Appendix F.1. As observed from Table F.3, the performance of all methods degrades as heterogeneity increases. FSLoRA consistently achieves higher accuracy.
Table F.3: Accuracy comparison under different data heterogeneity settings. FSLoRA maintains its advantage over the baselines as the data heterogeneity increases. The number of clients is set to $50$.
@@ F.3 Varying the Level of Data Heterogeneity
Appendix G Experiments on LLaMA-7B
Although our primary focus is on models suitable for client-side deployment, such as RoBERTa and LLaMA-3.2-3B models, we also include experiments on the larger LLaMA-7B model to demonstrate the scalability of FSLoRA in more complex models. Specifically, we fine-tune the LLaMA-7B model on the Wizard, Dolly-15k, and Alpaca datasets and evaluate it on $1444$ MMLU samples (available at: https://github.com/ATP-1010/FederatedLLM). For Wizard and Dolly-15k, we adopt the same heterogeneous data partitioning as [42]. Since the Alpaca dataset lacks a clear task or domain structure, we apply a uniform random partitioning strategy to distribute the data across clients. We tune the q_proj and v_proj modules and set the local LoRA ranks $k_{i}\!=\![64,32,16,16,8,8,4,4,4,4]$ for $10$ clients. The parameter settings are aligned with those in [42].
Table G.1: Performance comparison on LLaMA-7B model.
@@ Appendix G Experiments on LLaMA-7B
As shown in Table G.1, FSLoRA achieves the highest average performance across all three datasets compared to baselines. These results demonstrate FSLoRA’s potential for effective fine-tuning with the large-scale LLaMA-7B model under heterogeneous client settings.
Appendix H Further Experiments
@@
In this section, we provide additional results, including detailed per-task comparisons from the commonsense reasoning benchmark corresponding to Figures LABEL:fig:rank_varying and LABEL:fig:global_rank. In addition, we further investigate the impact of the number of local updates $H$ on the convergence, the robustness of FSLoRA under dynamic sketching ratio, and the integration of communication compression and sketching.
H.1 Further Details on the Ablation Study
@@
Impact of sketching: In Figure 6, we compare the performance of FSLoRA with and without sketching on eight tasks from the commonsense reasoning benchmark using the LLaMA-3.2-3B model. Notably, FSLoRA without sketching is equivalent to the vanilla Federated LoRA. For FSLoRA with sketching, we apply a uniform sketching ratio of $k_{i}/r=0.5$ across all distributed clients. The uploading budget for each client is set to $200$ times the size of the full global LoRA modules at the corresponding rank. It is clear that FSLoRA with sketching consistently outperforms its non-sketched counterpart across these eight tasks, demonstrating the effectiveness of sketching in improving performance.
@@
Figure 6: Comparison of FSLoRA with and without sketching, with an upload budget $200\times$ the global LoRA module size at each rank. This is based on the commonsense reasoning benchmark and the LLaMA-3.2-3B model. We observe that the sketching mechanism improves performance across all considered tasks. The average accuracy of the eight tasks is shown in Figure LABEL:fig:rank_varying.
@@
Figure 7: Impact of the rank of global LoRA modules on FSLoRA, given a fixed rank for the updated submatrices at the clients. This is based on the commonsense reasoning benchmark and the LLaMA-3.2-3B model. Overall, FSLoRA demonstrates improved performance as the global rank increases. The average accuracy of the eight tasks is shown in Figure LABEL:fig:global_rank.
@@
Impact of the global rank: In Figure 7, we present the impact of the rank of the global LoRA modules on FSLoRA’s performance across eight tasks from the commonsense reasoning benchmark. We consider four configurations: 1) $r=8,~k_{i}/r=1$, 2) $r=16,~k_{i}/r=0.5$, 3) $r=32,~k_{i}/r=0.25$, and 4) $r=64,~k_{i}/r=0.125$. The rank of submatrices updated by the clients at each iteration remains consistent across all configurations (i.e., $k_{i}=8$), ensuring that the communication and computational resources on the client side are kept fixed for all cases. In the ARC-Easy task, performance decreases as the rank increases to $64$, potentially due to overfitting. In general, FSLoRA shows improved performance as the rank increases.
H.2 Impact of Local Updates
Based on the commonsense reasoning benchmark and the LLaMA-3.2-3B model, we evaluated the convergence behavior of FSLoRA under varying numbers of local updates (i.e., $H$). The experimental results are presented in Figure 8. In the low-to-moderate regime of local updates (i.e., $H\in{10,20,100}$), FSLoRA demonstrates a clear acceleration in convergence as $H$ increases. For example, moving from $H=10$ to $H=20$ substantially reduces the number of communication rounds required to reach the same level of testing accuracy, while further increasing $H$ to $100$ yields even faster progress toward convergence. This observation indicates that a moderate increase in local updates allows clients to improve communication efficiency. However, when the number of local updates is pushed beyond this range (e.g., $H=200$), no additional convergence gain is observed. These findings align well with our theoretical analysis in Section 4, which shows that FSLoRA can achieve a convergence speedup under certain conditions on $H$.
Figure 8: Impact of the number of local updates on FSLoRA’s convergence. This is based on the commonsense reasoning benchmark and the LLaMA-3.2-3B model. In a certain range, i.e., from 10 to 100, FSLoRA achieves a fast convergence as $H$ increases.
H.3 Dynamic Sketching Ratios
While our primary focus is on developing a heterogeneous federated LoRA method under a standard static setup, following prior works [42, 8, 1], the proposed FSLoRA algorithm can be naturally extended to dynamic, time-varying resource environments. The modification is straightforward: we allow the sparsity levels, corresponding to the sketching ratios of FSLoRA, of the matrices in the sketching set $\mathcal{S}_{i}$ in Algorithm 1 to become time-varying, while keeping the remaining steps unchanged.
We empirically validate the effectiveness of FLoRA under this dynamic setting. In the simulation, we group clients into three capability levels low, medium, and high, assigned sketching ratio ranges $[0.125,0.25]$, $[0.25,0.5]$, and $[0.5,1.0]$, respectively, to balance local training latencies across groups. Within each range, the sketching ratios are allowed to vary dynamically. The results, reported in Table H.1, show that FSLoRA maintains comparable performance when moving from the static to the dynamic case, demonstrating its robustness under time-varying sketching ratios.
Table H.1: The performance of FSLoRA under static and dynamic sketching ratios. This is based on the commonsense reasoning benchmark and the LLaMA-3.2-3B model. FSLoRA maintains comparable performance when moving from the static to the dynamic case.
Ratios ARC-c ARC-e BoolQ HSwag OBQA PIQA SIQA Wino Avg.
Static 76.1 87.1 70.0 81.7 81.4 82.6 76.4 78.9 79.3
[…]
H.4 Integration of Sketching and Top-k Compression
@@
Building on the commonsense reasoning benchmark and the LLaMA-3.2-3B model, we further explore the integration of two orthogonal techniques, sketching and top-k compression, to further reduce the uplink communication overhead of clients in FSLoRA.
Specifically, with sketching, each client activates and updates submatrices of the global LoRA weights, $\mathbf{b}_{j}]_{j\in\mathcal{I}_{i}},[\mathbf{a}_{j}^{\top}]_{j\in\mathcal{I}_{i}}$, which are selected at the beginning of each round:
$\mathbf{BS}_{i}\mathbf{A}=\sum_{j\in\mathcal{I}_{i}}\frac{r}{k_{i}}\mathbf{b}_{j}\mathbf{a}_{j}^{\top}=\frac{r}{k_{i}}[\mathbf{b}_{j}]_{j\in\mathcal{I}_{i}}[\mathbf{a}_{j}^{\top}]_{j\in\mathcal{I}_{i}},$
where $\mathbf{b}_{j}$ and $\mathbf{a}_{j}^{\top}$ denote the $j$-th column of module $\mathbf{B}$ and the $j$-th row of module $\mathbf{A}$, respectively. By limiting updates to submatrices $\mathbf{b}_{j}]_{j\in\mathcal{I}_{i}}$ and $\mathbf{a}_{j}^{\top}]_{j\in\mathcal{I}_{i}}$, FSLoRA reduces communication and computation. To further reduce communication cost, we can apply Top-$k$ compression to the uploading stage. For instance, instead of sending the full $\Delta[\mathbf{b}_{j}]_{j\in\mathcal{I}_{i}}$, each client transmits the compressed update $\operatorname{Top}_{k}(\Delta[\mathbf{b}_{j}]_{j\in\mathcal{I}_{i}})$. Sketching selects the update submatrix at the beginning of each round, while compression further reduces its transmission cost at the uploading stage. These two techniques operate at different stages and can jointly improve communication efficiency.
In our setup, the compression ratio is fixed at $0.5$ for all methods, while the sketching ratio $k_{i}/r$ varies over $\{0.125,0.25,0.5,1\}$. Notably, FSLoRA with sketching ratio $k_{i}/r=1$ corresponds to the vanilla Federated LoRA (i.e., without sketching). Figure 9 plots testing accuracy versus communication overhead, where the x-axis represents the amount of data uploaded per client (in MB), assuming parameters are stored in float 32 precision. The results clearly show that integrating sketching with top-k compression further improves communication efficiency: methods with lower sketching ratios consistently achieve higher accuracy under the same communication budget, highlighting the potential of FSLoRA for scalable and communication-efficient collaborative LLM fine-tuning.
Appendix I Implementation Details
I.1 Details on Hyperparameters
Table I.1: The hyperparameters for RoBERTa & GLUE and LLaMA-3.2-3B & commonsense reasoning benchmarks.
Hyperparameter RoBERTa & GLUE LLaMA-3.2-3B & commonsense reasoning
Dirichlet parameter 0.1 0.1
Batch size 16 16
LoRA dropout rate 0.1 0.1
Learning rate, $\gamma$ 5e-4 3e-4
Communication round, $T$ 200 200
Local iteration number, $H$ 50 20
I.2 Details on Datasets
I.2.1 GLUE Benchmark
[…]
      Input: A single sentence.
[…]
      Input: A single sentence.
    • Output: A label indicating positive or negative sentiment.
[…]
      Input: Two sentences (‘sentence1’ and ‘sentence2’).
    • Output: A label indicating either equivalent or not equivalent.
[…]
      Input: Two questions.
    • Output: A label indicating duplicate or not duplicate.
[…]
    • Output: A label indicating entailment, contradiction, or neutral.
[…]
      Input: A question and a sentence.
  • RTE provides pairs of sentences to see if one implies the other.
[…]
      Input: Two sentences (‘sentence1’ and ‘sentence2’)
I.2.2 Commonsense Reasoning Benchmark
The input template, i.e., prompt format for these datasets is detailed in Table I.2.
Table I.2: The prompt template of the commonsense reasoning datasets [19].
Dataset Input Template
Appendix J Proof of the Theoretical Results
J.1 Preliminaries
Before presenting the proof of the main results, we first introduce some preliminary facts that will be used later. Throughout this work, $\|\cdot\|$ denotes the Frobenius norm when applied to a matrix and the $\ell_{2}$ norm when applied to a vector.
Lemma J.1.
@@ I.2.2 Commonsense Reasoning Benchmark
Lemma J.2.
[40, Lemma 2] Suppose a sequence of random matrices $\left\{{\bf P}_{i}\right\}_{i=1}^{N}$ satisfy $\mathbb{E}\left[{\bf P}_{i}\mid{\bf P}_{i-1},\!{\bf P}_{i-2},\ldots,{\bf P}_{1}\right]={\bf 0},\forall i.$ Then,
@@
Lemma J.3.
[23, Lemma 17] For any $a_{0}\geq 0,b\geq 0,c\geq 0,d>0$, there exist a constant $\eta\leq\frac{1}{d}$ such that
@@
Lemma J.4 (Random sketching bounds).
@@
This completes the proof of Lemma J.4. ∎
J.2 Proof of Lemma 3.2
@@
The proof of Proposition J.5 is deferred to Appendix J.4. With this proposition, we are ready to prove Theorem 4.4.
In FSLoRA, the update direction in (4) corresponds to the negative stochastic gradient of $\ell({\bf W}_{0}+{\bf B}{\bf S}{\bf A},\xi)$ with respect to ${\bf B};{\bf A}]$ for a given sketch ${\bf S}_{i}^{t}$. We have defined $\tilde{\ell}({\bf X},\xi;{\bf S})=\ell({\bf W}_{0}+{\bf B}{\bf S}{\bf A},\xi)$. The iterative equation for the proposed FSLoRA algorithm thus can be written as
@@
where ${\bf g}^{t,h}_{i}$ denotes the stochastic gradient $\nabla_{\bf X}\tilde{\ell}({\bf X}^{t,h}_{i},\xi_{i}^{t,h};{\bf S}_{i}^{t})$. Based on the smoothness of $f^{\mathcal{S}}(\bf X)$, i.e., Proposition J.5, we have
@@
where the last inequalities follow Jensen’s inequality and Proposition J.5.
@@
where the inequality follows the fact that $\mathbb{E}[\sum_{h=0}^{H-1}({\bf g}^{t,h}_{i}-\nabla_{\bf X}f_{i}^{\mathcal{S}}({\bf X}_{i}^{t,h}))]=0$ and the independence between clients.
@@
where the equality follows Lemma J.2 and the inequality follows Assumption 4.2. For $\left\|\nabla_{\bf X}f_{i}^{\mathcal{S}}({\bf X}_{i}^{t,h})\right\|^{2}$, utilizing Assumption 4.3 and Proposition J.5, we have
@@
where $T_{3}=\frac{1}{NH}\sum_{i=1}^{N}\frac{r^{2}}{k_{i}^{2}}\sum_{h=0}^{H-1}\mathbb{E}\left\|{\bf X}_{i}^{t,h}-{\bf X}^{t}\right\|^{2}$. Combining (12), (13), and (J.3) yields
@@
Plugging inequality (J.3) into inequality (17) yeilds
@@
Telescoping the above inequality from $t=0$ to T1T-1, we have
@@
where $f^{*}$ denotes the lower bound of $f^{\mathcal{S}}({\bf X})$ and $\widetilde{L}=\kappa L$.
Applying Lemma J.3 to the above inequality and letting $d=H$, it follows that there exists a learning rate $\gamma\leq\min\{\frac{N}{24\rho(c_{h}+1)H\bar{L}},\frac{1}{8\sqrt{\widetilde{L}L(\rho+1)(c_{h}+1)}H},\frac{1}{H}\}$ such that
@@
This completes the proof of Theorem 4.4.
J.4 Proof of Proposition J.5
i) For illustration, we need to recover $\bf X$ to $\bf B;{\bf A}]$ in this proof. According to the definition of $\tilde{f}_{i}({\bf X};{\bf S})$ and $f_{i}({\bf B},{\bf A})$, we have